OpenStudy (anonymous):

Which rule would apply to the expression w^2w^3? A. Addition rule B. Power rule C. Distributive rule D. Quotient to difference rule

OpenStudy (debbieg):

I'm assuming you mean which rule of exponents? You have \(\large w^2 w^3\) which is really just: \(\large (w\cdot w)(w\cdot w \cdot w)=w\cdot w\cdot w\cdot w\cdot w=??\) So, what do you need to do with the exponents? :)